OpenStudy (anonymous):

(6x/5y)x(10y/8x) simplify

OpenStudy (jdoe0001):

\(\bf \cfrac{6x}{5y}\times\cfrac{10y}{8x}\implies \cfrac{6x\times10y}{5y\times 8x}\)

OpenStudy (luigi0210):

Cross cancel: \[\Large \frac{6x}{5y}*\frac{10y}{8x}\]

OpenStudy (anonymous):

Cancel what?

OpenStudy (luigi0210):

If it helps, combine them like jdoe did, then separate them again like: \[\Large \frac{6x}{8x}*\frac{10y}{5y}\]

OpenStudy (luigi0210):

Now simplify

OpenStudy (anonymous):

both the denominator and numerator?

OpenStudy (luigi0210):

Yea, then combine them back together

OpenStudy (anonymous):

3/4 x 5y/5y?

OpenStudy (luigi0210):

3/4*2, yes So the final answer is..

OpenStudy (anonymous):

6/4?

OpenStudy (luigi0210):

Simplify it down one more time

OpenStudy (anonymous):

is it a decimal?

OpenStudy (luigi0210):

Take a 2 out of them?

OpenStudy (anonymous):

Out of where?

OpenStudy (anonymous):

I divided 6 and 4

OpenStudy (luigi0210):

divide 6 and 4 by 2..

OpenStudy (anonymous):

3/2

OpenStudy (luigi0210):

There you go!

OpenStudy (anonymous):

Im mad slow

OpenStudy (anonymous):

Thanks man

OpenStudy (luigi0210):

Anytime!

OpenStudy (shamil98):

60xy/40xy = 6/4 = 3/2